# Global minimality of the Hopf map in the Faddeev–Skyrme model with large coupling constant

### v1

@@ Abstract.
We prove that, modulo rigid motions, the Hopf map is the unique minimizer of the Faddeev–Skyrme energy in its homotopy class, for a sufficiently large coupling constant.
@@ 1. Introduction
The Skyrme model, introduced in [29], is a classical nonlinear $O(3)$-sigma model which has proved successful in quantum field theory. The static fields in Skyrme’s model, which are called Skyrmions, can be mathematically described as critical points of the energy
@@
where $\mathbb{M}^{3}$ is a Riemannian manifold [^1] Here, $\wedge$ is the natural wedge product on $T^{*}\mathbb{M}^{3}\otimes\mathbb{R}^{4}$ characterized by $(\alpha\otimes a)\wedge(\beta\otimes b)=(\alpha\wedge\beta)\otimes(a\wedge b)$.. Skyrmions are a particular example of topological solitons, and we refer the reader to the monograph [25] for further physical background. The existence of maps minimizing $\mathcal{S}$ in suitable homotopy classes has been proved rigorously in the works [8, 9], in the case $\mathbb{M}^{3}=\mathbb{R}^{3}$. In this case, finite-energy maps exhibit rapid decay at infinity to a constant value $b\in{\mathbb{S}}^{3}$, so that they can be considered as maps from $\mathbb{R}^{3}\cup\{\infty\}\cong{\mathbb{S}}^{3}\to{\mathbb{S}}^{3}$.
A further refinement of Skyrme’s model was proposed by Faddeev [11], considering Skyrme’s model in the case where the maps take values in an equator $\mathbb{S}^{2}\subset\mathbb{S}^{3}$. As in Skyrme’s model, there is a satisfactory existence theory for this problem, especially in the case of compact $\mathbb{M}^{3}$ [1, 22, 23, 33], the case of non-compact $\mathbb{M}^{3}$ being more challenging [22, 23]. In the simplest non-compact example $\mathbb{M}^{3}=\mathbb{R}^{3}$, and following Ward [33], one can consider an approximation of $\mathbb{R}^{3}$ by $\mathbb{S}^{3}_{\rho}$, the sphere of radius $\rho\gg 1$. Thus, considering the energy on ${\mathbb{S}}^{3}_{\rho}$, after a change of variables and normalization by a factor $\rho^{-1}$, one is led to a perturbed version of the Faddeev–Skyrme energy, namely
@@
Homotopy classes of smooth maps from $\mathbb{S}^{3}$ to ${\mathbb{S}}^{2}$, corresponding to elements of $\pi_{3}(\mathbb{S}^{2})\cong\mathbb{Z}$ are classified by their Hopf invariant (or Hopf charge in the physics literature) [34], which is defined for $u\in C^{\infty}({\mathbb{S}}^{3};{\mathbb{S}}^{2})$ as
@@
where $\beta$ is any 1-form on ${\mathbb{S}}^{3}$ such that $d\beta=u^{\ast}\omega_{{\mathbb{S}}^{2}}$. This definition can be meaningfully extended to the class of finite Faddeev–Skyrme energy maps, see [2] or Section 3 below.
A basic question concerning the above models is whether one can characterize minimizers in some simple cases; the non-convexity of the problems makes this a rather difficult question. For the Skyrme model with $\mathbb{M}^{3}=\mathbb{S}^{3}_{\rho}$, it is conjectured [24] that the homothety $x\mapsto x/\rho$ is a minimizer in its homotopy class whenever $0<\rho\leq\sqrt{2}$; this map is known to be unstable whenever $\rho>\sqrt{2}$. At present minimality is only known for $0<\rho\leq\sqrt{3/2}<\sqrt{2}$, see [28] and the references therein.
@@
$\pi\colon\mathbb{C}\cup\{\infty\}\to{\mathbb{S}}^{2}\,,\quad\pi(z):=\left(\frac{2z}{|z|^{2}+1},\frac{|z|^{2}-1}{|z|^{2}+1}\right)\,,$ (1.4)
@@
For the Faddeev–Skyrme model it is also known [33] that $h$ is unstable for $\rho>\sqrt{2}$, while it is linearly stable for $0<\rho\leq\sqrt{2}$, cf. [17, 30], similarly to the homothety in the Skyrme model. It is therefore conjectured that, in the latter case, $h$ is a minimizer of the energy in its homotopy class.
Actually, as pointed out [^2] Specifically, cf. the paragraph below Theorem 1.4 in [31]. in [31], the above conjectured minimality property of the Hopf map was unknown for any value of the coupling constant $0<\rho\leq\sqrt{2}$. In this paper we give a positive answer to the above conjecture, for sufficiently small values of $\rho$:
@@
There exists $\rho_{0}\in(0,\sqrt{2})$ such that, for every $0<\rho\leq\rho_{0}$ and $u\in\mathcal{U}_{\mathcal{FS}}$ with Hopf invariant equal to 1, it holds that
$\mathcal{FS}_{\rho}(u)\geq\mathcal{FS}_{\rho}(h)\,.$ (1.6)
Equality in (1.6) holds if and only if $u=h\circ R$ for some $R\in\textup{SO}(4)$.
The case where the energy consists only of the second term in (1.1), which formally corresponds to the limit $\rho\to 0$, was previously considered in [31], where a short proof of the minimality of the Hopf map was provided.
It is also interesting to compare our results with another conjecture, this the Hopf map should minimize, in its homotopy class, the conformally-invariant energy
@@
Linear stability of $h$ in this case was already proved in [27], as well as minimality if the power 3 is replaced with 4, but otherwise this conjecture remains open; see also [7] for related results. On the one hand, there are several similarities between (1.7) and the Faddeev–Skyrme problem, cf. for instance [22, 23], and as we will see in this paper the required spectral analysis is also quite similar. On the other hand, the problems are at the same time fairly different, since (1.1) is not conformally invariant and the corresponding integrand is non-convex.
Our strategy to prove Theorem 1.1 elaborates on two already mentioned known facts about the Hopf map:
@@
    it is linearly stable for $0<\rho\leq\sqrt{2}$, cf. [17, 30];
@@
    it is minimal for the second term in the Faddeev-Skyrme energy (1.1), cf.  [31].
We establish quantitative versions of these two facts: loosely stated,
@@
    the Hopf map is (modulo symmetries) a strict local minimizer for $0<\rho<\sqrt{2}$;
@@
    if a map is almost minimal for the second energy term, then it must be close to the Hopf map (modulo symmetries).
Since the second energy term in (1.1) is dominant for small $\rho$, these strengthened assertions provide a proof of Theorem 1.1 : minimizers of the full energy must be almost minimal for the second energy term, hence close to the Hopf map thanks to (b’), and therefore equal to the Hopf map (modulo rotational symmetries) thanks to (a’). Here it is of course crucial that the closeness to the Hopf map provided by (b’) implies the closeness required by (a’). In fact, a version of (a’) established in [17] requires closeness in a too restrictive distance and cannot be combined with (b’). Another difficulty is related to the symmetries in these statements: the second energy term in (1.1), which depends only on the closed 2-form $\alpha=u^{*}\omega_{{\mathbb{S}}^{2}}$, has more symmetries than the full energy. Hence, the closeness modulo symmetries provided by (b’) cannot imply the closeness modulo symmetries required by (a’). But this difficulty is resolved thanks to an idea introduced by Rivière [27, 28] and used also in [17], which allows to apply the outlined strategy using a relaxed energy on closed $2$-forms.
Plan of the paper: In Section 2 we gather some notation and basic properties related to the Hopf map and the Hodge Laplacian on $\mathbb{S}^{3}$. In Section 3 we prove several useful facts about the Hopf invariant for maps from ${\mathbb{S}}^{3}$ to ${\mathbb{S}}^{2}$ with finite Faddeev–Skyrme energy, and include a concise proof of its integrality in this lower-regularity case. In Section 4, we establish local coercivity properties for a relaxed energy functional on closed $2$-forms. In Section 5, we turn this local coercivity estimate into a global stability estimate for the relaxed energy, for sufficiently large values of the coupling constant $\rho^{-2}$, and with this as our main ingredient, we conclude the proof of Theorem 1.1.
@@ 2.1. Some basic identities involving differential forms
This gives rise to an alternating vector-valued 2-form $du(x)\wedge du(x)\in{\bigwedge^{2}T_{x}^{\ast}{\mathbb{S}}^{3}\otimes\mathbb{R}^{3}}$, which in coordinates is expressed as
@@
where as usual $\widehat{\cdot}$ denotes that the corresponding term is ommited. Thus, when $n=2$, its pull-back by $u$ is
$\begin{split}u^{*}\omega_{{\mathbb{S}}^{2}}&=u^{1}du^{2}\wedge du^{3}+u^{2}du^{3}\wedge du^{1}+u^{3}du^{1}\wedge du^{2}\\ &=\frac{1}{2}u\cdot du\wedge du=\sum_{1\leq j<k\leq 3}u\cdot(\partial_{j}u\times\partial_{k}u)\,dx^{j}\wedge dx^{k}\,,\end{split}$ (2.3)
where $\cdot$ and $\times$ denote the Euclidean inner and outer product respectively. Since $|u|=1$, (2.1) and (2.3) imply that
@@
Throughout this paper we will use systematically Hodge theory for differential forms with coefficients in Sobolev spaces, and we refer the reader to [18, §5] for a comprehensive discussion of this theory. Here we record the following simple lemma:
@@
[…]
@@
    $\Big{|}\int_{\mathbb{S}^{3}}\varphi\wedge d\varphi\Big{|}\leq C\|d\varphi\|_{L^{2}({\mathbb{S}}^{3})}^{2}\,.$ (2.6)
@@
The claim in (i) follows by the formula $d(\alpha\wedge\beta)=d\alpha\wedge\beta-\alpha\wedge d\beta$ and Stokes’ Theorem. Regarding (ii), let $\overline{\varphi}:=\fint_{{\mathbb{S}}^{3}}\varphi\in\bigwedge^{1}T^{*}{\mathbb{S}}^{3}$ be the mean value of $\varphi$, intended component-wise.
Then,
@@
which the reader can find in [18, §4]. ∎
We will also use, often implicitly, the characterization of the De Rham cohomology groups of spheres. In particular, for $\mathbb{S}^{3}$ we have that
@@
cf. [21, Theorem 17.21]. In particular closed 1- and 2-forms on $\mathbb{S}^{3}$ are exact, and the same holds for forms with Sobolev coefficients, cf. [18, §5].
@@ 2.2. Spectral properties of the Hodge-Laplacian on closed forms
In fact, it is also useful to consider the square root of $\Delta$ which, restricted to closed 2-forms, becomes the operator $d*$. The corresponding eigenvalue equation is then
@@
Note that if $\alpha$ solves (2.8) then $\Delta\alpha=\lambda^{2}\alpha$. Since the spectrum of $\Delta$ on $\Omega^{2}(\mathbb{S}^{3})$ is the set $\{(k+2)^{2}\}_{k\in\mathbb{N}}$, cf. [16, 19], it follows that for (2.8) to be solvable we must have $\lambda=\pm(k+2)$. In fact (2.8) is solvable for all such values, see again [16, 19] as well as [27, Proposition IV.2], and we have the following precise description of the corresponding eigenspaces:
Proposition 2.2 (Spectrum of $d\,*$).
@@
are the restriction to $\mathbb{S}^{3}$ of the self-dual (respectively anti-self-dual) closed and co-closed 2-forms on $\mathbb{R}^{4}$ which are homogeneous and polynomial of degree $k$. These spaces induce an $L^{2}$-orthogonal decomposition
@@
We recall that a form $\alpha\in\Omega^{2}(\mathbb{R}^{4})$ is said to be self-dual/anti-self-dual if $*\alpha=\pm\alpha$.
@@
Since $**=\mathrm{Id}$ on $\Omega^{2}(\mathbb{R}^{4})$ and the Hodge-star operation is an isometry, the spaces of self-dual and anti-self-dual forms give a pointwise-orthogonal decomposition of $\Omega^{2}(\mathbb{R}^{4})$. In the case $k=0$, the elements in $E_{0}^{\pm}$ have a specially simple description, since they are the restrictions to $\mathbb{S}^{3}$ of constant-coefficient 2-forms on $\mathbb{R}^{4}$, and hence they can be identified with the self-dual (or anti-self-dual) constant-coefficient 2-forms on $\mathbb{R}^{4}$. For instance, we have
@@
with a similar basis for $E_{0}^{-}$, up to flipping the signs in the above basis. Note in particular that $\dim E_{0}^{\pm}=3$.
@@
$\forall\omega_{1},\omega_{2}\in E_{0}^{\pm}\text{ with }|\omega_{1}|=|\omega_{2}|\text{ there exists }\widetilde{R}\in SO(4)\text{ such that }\omega_{2}=\widetilde{R}^{*}\omega_{1}\,.$ (2.12)

### v2

@@ Abstract.
We prove that, modulo rigid motions, the Hopf map is the unique minimizer of the Faddeev–Skyrme energy in its homotopy class, provided that the radius of the target 2-sphere is not smaller than the radius of the domain 3-sphere.
@@ 1. Introduction
The Skyrme model, introduced in [skyrme1961non], is a classical nonlinear $O(3)$-sigma model which has proved successful in quantum field theory. The static fields in Skyrme’s model, which are called Skyrmions, can be mathematically described as critical points of the energy
@@
where $\mathbb{M}^{3}$ is a Riemannian manifold [^1] Here, $\wedge$ is the natural wedge product on $T^{*}\mathbb{M}^{3}\otimes\mathbb{R}^{4}$ characterized by $(\alpha\otimes a)\wedge(\beta\otimes b)=(\alpha\wedge\beta)\otimes(a\wedge b)$.. Skyrmions are a particular example of topological solitons, and we refer the reader to the monograph [Manton2004] for further physical background. The existence of maps minimizing $\mathcal{S}$ in suitable homotopy classes has been proved rigorously in the works [Esteban1986, Esteban2004], in the case $\mathbb{M}^{3}=\mathbb{R}^{3}$. In this case, finite-energy maps exhibit rapid decay at infinity to a constant value $b\in{\mathbb{S}}^{3}$, so that they can be considered as maps from $\mathbb{R}^{3}\cup\{\infty\}\cong{\mathbb{S}}^{3}\to{\mathbb{S}}^{3}$.
A further refinement of Skyrme’s model was proposed by Faddeev [faddeev1976some], considering Skyrme’s model in the case where the maps take values in an equator $\mathbb{S}^{2}\subset\mathbb{S}^{3}$. As in Skyrme’s model, there is a satisfactory existence theory for this problem, especially in the case of compact $\mathbb{M}^{3}$ [Auckly2005, lin2004existence, lin2007energy, ward1999], the case of non-compact $\mathbb{M}^{3}$ being more challenging [lin2004existence, lin2007energy]. In the simplest non-compact example $\mathbb{M}^{3}=\mathbb{R}^{3}$, and following Ward [ward1999], one can consider an approximation of $\mathbb{R}^{3}$ by $\mathbb{S}^{3}_{\rho}$, the sphere of radius $\rho\gg 1$. Thus, considering the energy on ${\mathbb{S}}^{3}_{\rho}$, after a change of variables and normalization by a factor $\rho^{-1}$, one is led to a perturbed version of the Faddeev–Skyrme energy, namely
@@
Homotopy classes of smooth maps from $\mathbb{S}^{3}$ to ${\mathbb{S}}^{2}$, corresponding to elements of $\pi_{3}(\mathbb{S}^{2})\cong\mathbb{Z}$ are classified by their Hopf invariant (or Hopf charge in the physics literature) [whitehead1947expression], which is defined for $u\in C^{\infty}({\mathbb{S}}^{3};{\mathbb{S}}^{2})$ as
@@
where $\beta$ is any 1-form on ${\mathbb{S}}^{3}$ such that $d\beta=u^{\ast}\omega_{{\mathbb{S}}^{2}}$. This definition can be meaningfully extended to the class of finite Faddeev–Skyrme energy maps, see [auckly2010pontrjagin] or Section 3 below.
A basic question concerning the above models is whether one can characterize minimizers in some simple cases; the non-convexity of the problems makes this a rather intriguing question. For the Skyrme model with $\mathbb{M}^{3}=\mathbb{S}^{3}_{\rho}$, it is conjectured [Manton1986] that the homothety $x\mapsto x/\rho$ is a minimizer in its homotopy class whenever $0<\rho\leq\sqrt{2}$; this map is known to be unstable whenever $\rho>\sqrt{2}$. At present minimality and uniqueness modulo isometries is only known for $0<\rho\leq\sqrt{3/2}<\sqrt{2}$, see [riviere1998remark] and the references therein.
@@
$\pi\colon\mathbb{C}\cup\{\infty\}\to{\mathbb{S}}^{2}\,,\quad\pi(z):=\left(\frac{2z}{|z|^{2}+1},\frac{|z|^{2}-1}{|z|^{2}+1}\right)\,$,
@@
For the Faddeev–Skyrme model it is also known [ward1999] that $h$ is unstable for $\rho>\sqrt{2}$, while it is linearly stable for $0<\rho\leq\sqrt{2}$, cf. [isobe2008, speight2007], similarly to the homothety in the Skyrme model. It is therefore conjectured that, in the latter case, $h$ is the unique (modulo isometries) minimizer of the energy in its homotopy class.
Actually, as pointed out [^2] Specifically, cf. the paragraph below Theorem 1.4 in [speight2011]. in [speight2011], the above conjectured minimality property of the Hopf map was unknown for any value of the coupling constant $0<\rho\leq\sqrt{2}$. In this paper we give a positive answer to the above conjecture for a certain range of $\rho$:
@@
[…]
$\mathcal{FS}_{\rho}(u)\geq\mathcal{FS}_{\rho}(h)\,$,
with equality if and only if $u=h\circ R$ for some $R\in\textup{SO}(4)$.
The case where the energy consists only of the second term in (1.1), which formally corresponds to the limit $\rho\to 0$, was previously considered in [speight2011], where a short proof of the minimality of the Hopf map was provided.
It is also interesting to compare our results with another natural conjecture, tha the Hopf map should minimize, in its homotopy class, the conformally-invariant energy
@@
Linear stability of $h$ in this case was already proved in [riviere1998], as well as minimality if the power 3 is replaced with 4, but otherwise this conjecture remains open; see also [CoronGulliver] for related results. On the one hand, there are several similarities between (1.5) and the Faddeev–Skyrme problem, cf. for instance [lin2004existence, lin2007energy], and as we will see in this paper the required spectral analysis is also quite similar. On the other hand, the problems are at the same time fairly different, since (1.1) is not conformally invariant and the corresponding integrand is non-convex.
A natural strategy to prove Theorem 1.1 would elaborate on two already mentioned known facts about the Hopf map:
@@
    it is linearly stable for $0<\rho\leq\sqrt{2}$, cf. [isobe2008, speight2007];
@@
    it is minimal for the second term in the Faddeev-Skyrme energy (1.1), cf. [speight2011].
If one could establish quantitative versions of these two facts: loosely stated that,
@@
    the Hopf map is (modulo symmetries) a strict $W^{1,2}$-local minimizer for $0<\rho<\sqrt{2}$;
@@
    if a map is almost minimal for the second energy term, then it must be close to the Hopf map (modulo symmetries),
then since the second energy term in (1.1) is dominant for small $\rho$, these strengthened assertions would provide a proof of Theorem 1.1 for a sufficiently small, though non-explicit range of $\rho$. Our proof is actually a more direct variant of the aforementioned strategy and it yields an explicit range. A key difficulty is related to the symmetries in these statements: the second energy term in (1.1), which depends only on the closed 2-form $\alpha=u^{*}\omega_{{\mathbb{S}}^{2}}$, has more symmetries than the full energy. But this difficulty is resolved thanks to an idea introduced by Rivière [riviere1998, riviere1998remark] and used also in [isobe2008], which allows to apply the outlined strategy using a relaxed energy on closed $2$-forms. Note that, in contrast to our setting, in [riviere1998remark] the relaxed energy is quadratic, cf. (1.10) therein, and thus the associated relaxed minimization problem can be solved explicitely.
Plan of the paper: In Section 2 we gather some notation and basic properties related to the Hopf map and the Hodge Laplacian on $\mathbb{S}^{3}$. In Section 3 we prove several useful facts about the Hopf invariant for maps from ${\mathbb{S}}^{3}$ to ${\mathbb{S}}^{2}$ with finite Faddeev–Skyrme energy, and include a concise proof of its integrality in this lower-regularity case. In Section 4, we establish global minimality properties for a relaxed energy functional on closed $2$-forms, the main result therein being Proposition 4.2. With this as our key ingredient, at the end of the section we obtain the minimality of the Hopf map as an immediate consequence, and in Section 5 we conclude the proof of Theorem 1.1 by verifying the uniqueness (modulo rotational symmetries) in the statement.
@@ 2.1. Some basic identities involving differential forms
[…]
@@
where as usual $\widehat{\cdot}$ denotes that the corresponding term is omitted. Thus, when $n=2$, its pull-back by $u$ is
[…]
where $\cdot$ and $\times$ denote the Euclidean inner and outer products respectively. Since $|u|=1$, (2.1) and (2.3) imply that
@@
Throughout this paper we will use systematically Hodge theory for differential forms with coefficients in Sobolev spaces, and we refer the reader to [iwaniec1999nonlinear, §5] for a comprehensive discussion of this theory. Here we record the following simple lemma.
@@
    $\int_{\mathbb{S}^{3}}\alpha\wedge d\beta=\int_{\mathbb{S}^{3}}d\alpha\wedge\beta\,.$ (2.5)
@@
    $\Big|\int_{\mathbb{S}^{3}}\varphi\wedge d\varphi\Big|\leq C\int_{{\mathbb{S}}^{3}}|d\varphi|^{2}\,.$ (2.6)
@@
The proof of (i) is immediate by the formula $d(\alpha\wedge\beta)=d\alpha\wedge\beta-\alpha\wedge d\beta$ and Stokes’ theorem. Regarding (ii), let $\overline{\varphi}:=\fint_{{\mathbb{S}}^{3}}\varphi\in\bigwedge^{1}T^{*}{\mathbb{S}}^{3}$ be the mean value of $\varphi$, intended component-wise. Then
@@
which the reader can find in [iwaniec1999nonlinear, §4]. ∎
Actually, in the proof of Proposition 4.3, cf. (4.10) therein, we will calculate the optimal value of the constant $C>0$ in (2.6). We will also use, often implicitly, the characterization of the De Rham cohomology groups of spheres. In particular, for $\mathbb{S}^{3}$ we have that
@@
cf. [lee2003smooth, Theorem 17.21]. In particular closed 1- and 2-forms on $\mathbb{S}^{3}$ are exact, and the same holds for forms with Sobolev coefficients, cf. [iwaniec1999nonlinear, §5].
@@ 2.2. Spectral properties of the Hodge-Laplacian on closed forms
In fact, it is also useful to consider the square root of $\Delta$ which, restricted to closed 2-forms, becomes the operator dd*. The corresponding eigenvalue equation is then
@@
Note that if $\alpha$ solves (2.8) then $\Delta\alpha=\lambda^{2}\alpha$. Since the spectrum of $\Delta$ on $\Omega^{2}(\mathbb{S}^{3})$ is the set $\{(k+2)^{2}\}_{k\in\mathbb{N}}$, cf. [ikeda1978, Iwasaki1979], it follows that for (2.8) to be solvable we must have $\lambda=\pm(k+2)$. In fact (2.8) is solvable for all such values, see again [ikeda1978, Iwasaki1979] as well as [riviere1998, Proposition IV.2], and we have the following precise description of the corresponding eigenspaces:
Proposition 2.2 (Spectrum of dd\,*).
@@
are the restrictions to $\mathbb{S}^{3}$ of the self-dual (respectively anti-self-dual) closed and co-closed 2-forms on $\mathbb{R}^{4}$ which are homogeneous and polynomial of degree $k$. These spaces induce an $L^{2}$-orthogonal decomposition
@@
We recall that a form $\alpha\in\Omega^{2}(\mathbb{R}^{4})$ is said to be self-dual/anti-self-dual if $\alpha=\pm\alpha$.
@@
Since $\mathrm{Id}$ on $\Omega^{2}(\mathbb{R}^{4})$ and the Hodge-star operation is an isometry, the spaces of self-dual and anti-self-dual forms give a pointwise-orthogonal decomposition of $\Omega^{2}(\mathbb{R}^{4})$. In the case $k=0$, the elements in $E_{0}^{\pm}$ have a specially simple description, since they are the restrictions to $\mathbb{S}^{3}$ of constant-coefficient 2-forms on $\mathbb{R}^{4}$, and hence they can be identified with the self-dual (or anti-self-dual) constant-coefficient 2-forms on $\mathbb{R}^{4}$. For instance, we have
@@
with a similar basis for $E_{0}^{-}$, up to flipping the middle signs in the above basis. Note in particular that $\dim E_{0}^{\pm}=3$.
@@
[…]